George Washington was born eleven years before Thomas Jefferson. In 1770 Washington's age was three years more than seven times the age of Jefferson in 1748. How old was each man in 1750?
Answer by ankor@dixie-net.com(12695)

:
Let G = G.W. age in 1750
Let T = T.J. age in 1750
:
George Washington was born eleven years before Thomas Jefferson.
T = G - 11
:
In 1770 Washington's age was three years more than seven times the age of Jefferson in 1748.
:
G + 20 = 7(T - 2) + 3
G + 20 = 7T - 14 + 3
G + 20 = 7T - 11
G = 7T - 11 - 20
G = 7T - 31
:
How old was each man in 1750?
:
Substitute (G-11) for T:
G = 7(G-11) - 31
G = 7G - 77 - 31
G = 7G - 108
G - 7G = - 108
-6G = -108
G = -108/-6
G = +18 yrs old G in 1750
Then
T = 7 yrs old in 1750
:
I checked out the actual dates of birth and it checks out:
G.W. born in 1732 and T.J was both in 1743